# VerifiableFL: Verifiable Claims for Federated Learning using Exclaves

### v3

@@ 5. Auditing of Verifiable Claims
To verify that the data transformations led to the final trained FL model, VerifiableFL provides a mechanism to audit EDRs. For this, it constructs a graph based on all EDRs (§5.1). Using the graph, auditor verifies claims about the FL job (§5.2).
@@ 5.1. Constructing an exclave dataflow graph
The EDRs from all participants (as collected by the EDR storage service) are used to construct a directed acyclic graph, which we refer to as the exclave dataflow graph (EDG). Fig. 4 shows a sample EDG: each vertex $v\in V$ is an EDR, which is defined by the tuple of $(\mathit{in},\mathit{code},\mathit{out})$ that refers to the data input, code and data output hashes from that exclave execution. Each edge $e\in E$ represents the (reversed) data flow between exclaves; it points from the downstream (consumer) vertex $v_{y}$ to the upstream (producer) vertex $v_{x}$:
(1) $e=\{(v_{x},v_{y})\mid v_{x},v_{y}\in V\text{ and }v_{x}.\mathit{out}\cap v_{y}.\mathit{in}\neq\emptyset\}$
The edge label represents the specific data that flows between the two exclave executions. It is defined as the non-empty intersection of the outputs of $v_{x}$ and the inputs of $v_{y}$ (since exclaves may have multiple inputs and outputs for a single execution):
(2) $l(v_{x},v_{y})=\{v_{x}.out\cap v_{y}.in\}$
We define the $\mathit{EDG}=(V,E,l)$ as a collection of vertices $V$, reversed dataflow edges $E$, and an edge label function $l$. Alg. 1 shows how VerifiableFL constructs an EDG from a collection of EDRs $v$ that are attested as $\sigma(v)$. For each EDR, VerifiableFL first verifies the attestation with the hardware manufacturer, e.g., using AMD’s Key Distribution System (Kaplan, 2023) (line 1). If the verification succeeds, VerifiableFL adds the EDR to the EDG as a vertex; if it fails, e.g., due to an invalid signature or other tampering, the vertex is ignored (see Fig. 5(a)).
VerifiableFL then adds edges by iterating over all pairs of vertices. [^1] In practice, the structure of the FL job can be used to ascertain which vertices should be connected, avoiding a naive iteration in Alg. 1. If the inputs of an EDR vertex match the outputs of another, an edge is added between the two vertices, labeled with the dataflow between the exclaves (lines 68). If the data was modified outside the exclave, e.g., by an adversary during transit, there is no match and no edge is added due to the inputs/outputs mismatch (see Fig. 5(b)).
Algorithm 1 Constructing an EDG
@@
2:  $EDG\leftarrow\{V\}$
3:  for each $v_{y}\in V$ do
4:    for each $\{v_{x}\mid v_{x}\in V\text{ and }v_{x}\neq v_{y}\}$ do
5:      if $v_{x}.\mathit{out}\cap v_{y}.in\neq\emptyset$ do
6:       $e\leftarrow(v_{x},v_{y})$
7:       $l(e)\leftarrow v_{x}.\mathit{out}\cap v_{y}.\mathit{in}$
8:       $EDG\leftarrow EDG\cup\{e,l(e)\}$
@@ 5.2. Checking verifiable claims
The constructed EDG is audited to verify claims about the final FL model in terms of the faithful execution of the FL training protocol. Next, we describe possible verifiable claims using examples of specific training-time deviations to illustrate how an auditor detects deviations. The auditor can be an FL participant, a consumer of the final FL model who needs to verify correct training, or a trusted regulator who audits the training process for compliance.
@@
Claim 4: Integrity of training dataset. A malicious data provider could add, remove, or modify its dataset after training has commenced (e.g., performing an adaptive poisoning attack (Zhang et al., 2023)). If the data is modified while being read by an exclave, the verification against the Merkle tree fails; if the dataset is modified between training rounds (see Fig. 5(d)), the hash of the modified dataset is recorded in an EDG, enabling the auditor to discover the dataset change.
@@ 6. VerifiableFL Implementation
In this section, we describe the VerifiableFL implementation (§6.1) and its main components (§6.2), as well as our implementation of runtime attestation (§6.3) and integrity-protected storage (§6.4).
@@ 6.1. Overview
A goal of the VerifiableFL implementation is to be FL framework-agnostic and support existing FL frameworks with minimal code changes. To achieve this, VerifiableFL adopts an orchestrator/worker model in which an untrusted FL framework orchestrates the FL job and tasks execution inside exclaves.
Fig. 6 shows how VerifiableFL integrates with an FL framework. VerifiableFL executes an FL training job as follows: the model provider submits an FL job description to NVFlare  1 . The FL framework orchestrates the job according to its description, and allocates FL tasks to the associated exclaves via Exclave Connectors  2 , which are VerifiableFL components responsible for creating and invoking exclaves. When an exclave receives a task, an Exclave Manager  3 forwards the task payload (e.g., global model) to a Task Executor. During task execution  4 , the task can access persistent storage via a Secure Storage module  5 , which provides integrity protection for on-disk datasets. When the task completes, a Runtime Attestor  6 attests the EDR of the exclave, and returns the execution result (e.g., local model) to the FL framework.
@@ 6.2. Components
In more detail, VerifiableFL has the following components:
FL Framework. VerifiableFL uses NVFlare (Roth et al., 2022), which is an open-source, distributed FL framework by NVIDIA. It provides FL task implementations for a wide range of FL algorithms, e.g., federated averaging, federated optimization, and secure aggregation, and other algorithmic building blocks, e.g., model trainers, model aggregators, and task filters. Tab. 1 shows a set of tasks used in a typical FL job. VerifiableFL executes all of these FL tasks inside exclaves.
Exclave Connectors are gRPC clients (Indrasiri and Kuruppu, 2020) that NVFlare uses to execute FL tasks in exclaves. They are implemented as NVFlare tasks, which make them compatible with the NVFlare API. To add an exclave, VerifiableFL only needs to include the corresponding Exclave Connector in NVFlare’s job description file. This design using Exclave Connector avoids modifications to the NVFlare framework code.
Exclaves. VerifiableFL implements exclaves using confidential VMs (CVMs) based on AMD SEV-SNP (Sev-Snp, 2020). Within the CVM, exclaves are implemented as gRPC servers that execute tasks from the FL framework.
Exclave Managers expose exclave interfaces to the NVFlare to support the orchestration of the FL job. Exclave Managers receive FL tasks from NVFlare, deserialize their payloads, and forward the payload to the Task Executor. They then take the execution results from the Task Executor and pass them to NVFlare.
@@
Runtime Attestor. The Runtime Attestor constructs EDRs and interacts with a secure processor for attestation. During task execution, it monitors the storage accesses and requests/responses passing through the Exclave Manager to capture exclaves’ input and output data. The Runtime Attestor records a hash of each data item together with metadata that describes both the intended usage (e.g., input or output) and properties (e.g., model or dataset). Once an input/output pair is collected, it constructs the EDR as a nested JSON object, in which the keys contain the usage and properties of the data, and the values contain data hashes. The Runtime Attestor invokes the secure processor’s signing API with a SHA-256 digest of the EDR. The secure processor signs it together with the system measurement, and returns the signed EDR.
6.3. Emulated attestation path
To avoid relying on software secrets, we emulate the attestation path using a virtual Trusted Platform Module (vTPM) as the secure processor in AMD SEV-SNP. The vTPM provides equivalent attestation primitives—including measurement and signing—and executes at the highest privilege level in the VM. Its signing key is endorsed by the PSP when the CVM launches. We use virtual machine privilege modules (VMPL) to create privileges in memory: the vTPM runs in VMPL0 (highest privilege), while the exclave runs in VMPL2 (lower privilege). The MMU and RMP enforces isolation so that the exclave in a lower privilege level cannot interfere with attestation process in a higher privilege. Attestation then uses an MMU-protected shared memory page, which is a VMPL2 page writable by both the vTPM and the exclave. For attestation, the vTPM reads the runtime claim’s hash from the shared page, extends it into the PCR23, signs the PCRs, and writes the signed report back.
@@ 6.4. Secure storage with commitment
The Secure Storage module provides integrity-preserving and verifiable storage for exclaves, which is used to protect both datasets and task code.
Datasets. VerifiableFL uses Linux’s dm-verity (kernel development community, 2024) for integrity protection, which checks accessed disk blocks against a Merkle tree. Data providers use the dm-verity tool to package datasets with a Merkle tree. The generated data commitment becomes a unique measurement over the dataset, which is included in EDRs by the Runtime Attestor. When setting up an exclave, VerifiableFL mounts the data disk and checks the expected data commitment against the Merkle tree.
@@ 7. Evaluation
We evaluate the performance of VerifiableFL to answer the following questions: (i) what is VerifiableFL’s end-to-end training performance compared to existing solutions? (§7.2) (ii) how does VerifiableFL scale with an increased number of data providers and GPUs? (§7.3) (iii) what is the performance breakdown of VerifiableFL? (§7.4) (iv) how much overhead does an additional exclave introduce? (§7.5) (v) what overhead does VerifiableFL’s runtime attestation introduce? (§7.6) (vi) does secure storage introduce read amplification in terms of throughput and latency? (§7.7)
@@ 7.1. Experimental setup
Testbed. We deploy confidential VMs (CVMs) on Azure. All CVMs run on AMD EPYC 7763v CPUs with SEV-SNP and vTPM enabled; CVMs with GPUs have one NVIDIA H100 GPU with Hopper Confidential Computing extensions (Rob Nertney, 2023) enabled. Specifically, we deploy 5 CVMs (EC4as_v5) to execute 1 model update exclave and 4 NVFlare clients; 5 CVMs (EC8as_v5) to execute 4 DP exclaves and 1 NVFlare server; and 1 CVM (EC16as_v5) as the aggregation exclave. Training exclaves are deployed on 4 CVMs with CPUs (DC32as_v5) or 4 CVMs with GPUs (NCC40ads_H100_v5). Each training exclave mounts an Azure P30 data disk, which has up to 200 MB/s of throughput and support 5,000 IOPS.
VerifiableFL implementation. We implement VerifiableFL in Python, C++, and Golang. It uses PyTorch v2.5.0 with CUDA 12.4 for the FL tasks, Azure Guest Attestation Library v1.0.2 for CVM attestation, and NVFlare v2.2.1 (Roth et al., 2022) as the FL framework.
Baselines. We compare VerifiableFL with the following baselines: (i) NVFlare (No security) executes the FL workflow with the NVFlare server and clients in regular VMs. Each NVFlare node runs in a monolithic VM with SEV-SNP feature disabled; and (ii) NVFlare (CVM) (Guo et al., 2024) hosts each NVFlare client, as well as the server, in its own monolithic CVM. Note that for a fair comparison, both NVFlare (CVM) and VerifiableFL use the vTPM for attestation.
Datasets. We use 3 public datasets for VerifiableFL evaluation: (i) Wikitext103 (Merity et al., 2016) consists of raw text with over 100M tokens from Wikipedia articles; (ii) NCBI-disease (Doğan et al., 2014) is a disease text mining dataset with 793 fully annotated PubMed abstracts; and (iii) CIFAR10 (Krizhevsky et al., 2009) has annotated $32\times 32$ pixel images with 50,000 training and 10,000 testing samples.
ML models. We use 6 widely-used ML models, covering three different categories: (i) decoder-only transformers—we use GPT2 (Radford et al., 2019) with 137M parameters, and a distilled version of GPT2, DistilGPT2 (Sanh et al., 2019), with 88M parameters; (ii) encoder-only transformers—we employ BERT (Kenton and Toutanova, 2019) with 110M parameters, and MobileBERT (Sun et al., 2020) with 25M parameters; and (iii) convolutional neural networks (CNN)—we use LeNet (LeCun et al., 1998), a 5-layer CNN, and VGG9 (Simonyan and Zisserman, 2014), a 9-layer CNN.
@@
Figure 8. Throughput with different number of clients
@@ 7.2. End-to-end throughput
We evaluate the end-to-end performance of VerifiableFL in terms of throughput. We run VerifiableFL with 1 FL server and 4 FL clients. Each client performs model training with either a CPU or GPU CVM, depending on the workload. After each training round, FL clients use the sparse vector differential privacy technique (SVT-DP) (Lyu et al., 2016) to add DP noise to the locally trained model. VerifiableFL runs each FL task for 10 rounds with federated averaging (FedAvg) aggregation. We measure the performance of VerifiableFL for our three ML categories: (i) language modeling; (ii) named-entity recognition; and (iii) image classification. Fig. 7 shows the measured end-to-end throughput.
@@
Language modeling. In this experiment, each client has a training exclave that executes the language modeling task in a GPU CVM (40 vCPUs; 320 GB RAM; 1 NVIDIA H100 GPU). We train both GPT2 and DistilGPT2 models on the Wikitext dataset, which is grouped into sequences with a length of 1024 tokens after tokenization. Fig. 7(a) shows the training throughput of the language modeling task. With DistilGPT2, VerifiableFL has a $7\%$ overhead compared to the NVFlare (No security) baseline; NVFlare (CVM) has up to $2\%$ overhead. With GPT2, when compared to NVFlare (No security), VerifiableFL and NVFlare (CVM) incur less than $10\%$ and $4\%$ overhead in training throughput, respectively. We observe that, in both cases, VerifiableFL is slower than NVFlare (CVM): in VerifiableFL, exclaves introduce runtime overhead due to the runtime attestations; and the distributed deployment of exclaves introduces network overhead when (de)serializing the ML model and transmitting the serialized ML model.
@@
Named-entity recognition. Training exclaves execute in CPU CVMs (32 vCPUs; 128 GB memory). We use the MobileBERT and BERT models on the NCBI-disease dataset. Fig. 7(b) shows the training throughput of the named-entity recognition task. With MobileBERT, VerifiableFL and NVFlare (CVM) are $7\%$ and $6\%$ slower than NVFlare (No security), respectively. With BERT, VerifiableFL incurs a $9\%$ overhead compared to NVFlare (No security), and NVFlare (CVM) has a $4\%$ performance cost. We conclude that, when training MobileBERT, VerifiableFL has less performance impact compared to the BERT training. Since MobileBERT is $4.3\times$ smaller than BERT, it results in less network communication between exclaves in VerifiableFL.
Image classification. We also execute the image classification task with CPU CVMs and the same setup as before. Here, we use the LeNet and VGG9 models with the CIFAR10 dataset. Fig. 7(c) shows the performance results. For VGG9 training, VerifiableFL and NVFlare (CVM) incur $9\%$ and $4\%$ overhead, respectively, when compared to NVFlare (No security). VerifiableFL incurs a significant overhead for LeNet training: it is $29\%$ slower than NVFlare (No security). Although VerifiableFL has a low network overhead when transmitting small models, LeNet spends little time on local model training. Therefore, VerifiableFL’s latency overhead, although small, substantially contributes to the overall time, which impacts performance.
@@
For moderate and large models (i.e., VGG9, MobileBERT, BERT, DistilGPT2, and GPT2), VerifiableFL’s overhead is less than $10\%$ compared to NVFlare (No security), which is similar to NVFlare (CVM)’s overhead of $6\%$. Since local model training occupies most of the time in the FL workflow when executing FL tasks, VerifiableFL’s performance is insensitive to the extra latency of runtime attestation and its network overhead. In contrast, small models (i.e., LeNet) require little training time. Here, VerifiableFL’s performance becomes sensitive to the runtime and I/O costs, resulting in a performance degradation.

### v4

@@ 5. Auditing of Verifiable Claims
To verify that the data transformations led to the final trained FL model, VerifiableFL provides a mechanism to audit EDRs. For this, it constructs a graph based on all EDRs (§5.1). Using the graph, an auditor verifies claims about the FL job (§5.2).
@@ 5.1. Constructing an exclave dataflow graph
The EDRs from all participants (collected by the EDR storage service) are used to construct a directed acyclic graph, which we refer to as the exclave dataflow graph (EDG). Fig. 4 shows a sample EDG: each vertex $v\in V$ is an EDR, which is defined by the tuple of $(\mathit{in},\mathit{code},\mathit{out})$ that refers to the data input, code and data output hashes from that exclave execution. Each edge $e\in E$ represents the (reversed) data flow between exclaves; it points from the downstream (consumer) vertex $v_{y}$ to the upstream (producer) vertex $v_{x}$:
(1) $e=\{(v_{y},v_{x})\mid v_{y},v_{x}\in V\text{ and }v_{y}.\mathit{in}\cap v_{x}.\mathit{out}\neq\emptyset\}$
The edge label represents the data that flows between the two exclave executions. It is defined as the non-empty intersection of the outputs of $v_{x}$ and the inputs of $v_{y}$ (since exclaves may have multiple inputs and outputs for a single execution):
(2) $l(v_{y},v_{x})=\{v_{y}.in\cap v_{x}.out\}$
We define the $\mathit{EDG}=(V,E,l)$ as a collection of vertices $V$, reversed dataflow edges $E$, and an edge label function $l$. For auditing, Alg. 1 shows how VerifiableFL constructs an EDG from a collection of EDRs $v$ that are attested as $\sigma(v)$. For each EDR, VerifiableFL first verifies the attestation with the hardware manufacturer, e.g., using AMD’s Key Distribution System (amd_kds) (line 1). If the verification succeeds, VerifiableFL adds the EDR to the EDG as a vertex; if it fails, e.g., due to an invalid signature or other tampering, the vertex is ignored (see Fig. 5(a)).
VerifiableFL then adds edges by iterating over all vertices. [^1] In practice, the structure of the FL job can be used to ascertain which vertices should be connected, avoiding naive iteration. If the inputs of an EDR vertex match the outputs of another, an edge is added between the two vertices, labeled with the dataflow between the exclaves (lines 1013). If the data was modified outside the exclave, e.g., by an adversary during transit, there is no match and no edge is added due to the mismatch between inputs and outputs (see Fig. 5(b)).
Algorithm 1 Constructing an EDG for auditing
@@
2:  $EDG\leftarrow\{V,E=\emptyset,l=\emptyset\}$
3:  // Step 1: Index vertices by input
4:  for each $\{v_{y}\in V\}$ do
5:    for each $\{input\in v_{y}.\mathit{in}\}$ do
6:      append $v_{y}$ to $vertexOf[input]$
7:  // Step 2: Create edge when input and output match
8:  for each $\{v_{x}\in V\}$ do
9:    for each $\{output\in v_{x}.\mathit{out}\}$
10:      if $\{vertexOf[output]\neq\emptyset\}$
11:       for each $\{v_{y}\in vertexOf[output]\}$
12:        $e\leftarrow(v_{y},v_{x})$
13:        $l(e)\leftarrow v_{y}.\mathit{in}\cap v_{x}.\mathit{out}$
14:        $E\leftarrow E\cup\{e\}$, $l\leftarrow l\cup l(e)$
15:  return $(V,E,l)$
@@ 5.2. Checking verifiable claims
The constructed EDG is audited to verify claims about the trained FL model. Next, we describe possible verifiable claims using examples of specific training-time deviations and explain how an auditor detects these deviations:
@@
Claim 4: Integrity of training dataset. A malicious data provider could add, remove, or modify its dataset after training has commenced (e.g., performing an adaptive poisoning attack (a3fl)). If the data is modified while being read by an exclave, the verification against the Merkle tree fails; if the dataset is modified between training rounds (see Fig. 5(d)), the hash of the modified dataset is recorded in an EDG, enabling the auditor to discover the dataset change.
@@ 6. VerifiableFL Implementation
In this section, we describe the VerifiableFL implementation (§6.1) and its components (§6.2), as well as our implementation of exclave attestation (§6.3) and integrity-protected storage for data (§6.4).
@@ 6.1. Overview
A goal of the VerifiableFL implementation is to be compatible with different FL frameworks and to support existing FL frameworks with minimal code changes. To achieve this, VerifiableFL adopts an orchestrator/worker model in which an untrusted FL framework still orchestrates the FL job and tasks execution inside exclaves.
Fig. 6 shows how VerifiableFL integrates with an FL Framework: the model provider submits an FL job description to the FL Framework  1 . The FL Framework orchestrates the job according to its description, and allocates FL tasks to the associated exclaves via Exclave Connectors  2 , which are VerifiableFL components responsible for creating and invoking exclaves. When an exclave receives a task, an Exclave Manager  3 forwards the task payload (e.g., global model) to a Task Executor. During task execution  4 , the task can access persistent storage via a Secure Storage module  5 , which provides integrity protection for on-disk datasets. When the task completes, a Runtime Attestor  6 attests the EDR of the exclave, and returns the execution result (e.g., local model) to the FL Framework.
Table 1. FL tasks deployed in exclaves
@@ 6.2. Components
VerifiableFL has the following components:
FL Framework. VerifiableFL uses NVFlare (roth2022nvidia), which is an open-source, distributed FL framework by NVIDIA. It provides FL task implementations for many FL algorithms, e.g., federated averaging, federated optimization, and secure aggregation, and other algorithmic building blocks, e.g., model trainers, model aggregators, and task filters. Tab. 1 shows a set of tasks used in a typical FL job. VerifiableFL executes all of these FL tasks inside exclaves.
Exclave Connectors are gRPC clients (grpc) that NVFlare uses to execute FL tasks in exclaves. They are implemented as NVFlare tasks, which makes them compatible with the NVFlare API. To add an exclave, VerifiableFL includes the corresponding Exclave Connector in NVFlare’s job description file. This approach of using Exclave Connectors avoids modifications to the NVFlare code.
Exclaves. VerifiableFL implements exclaves as confidential VMs (CVMs) based on AMD SEV-SNP (sev2020strengthening). Within the CVM, exclaves are implemented as gRPC servers that execute FL tasks.
Exclave Managers expose exclave interfaces to NVFlare to support the orchestration of FL jobs. Exclave Managers receive FL tasks from NVFlare, deserialize their payloads, and forward the payload to the Task Executor. They then take the execution results from the Task Executor and pass them to NVFlare.
@@
Runtime Attestor. The Runtime Attestor constructs EDRs and interacts with a secure processor for attestation. During task execution, it monitors the storage accesses and requests/responses passing through the Exclave Manager to capture exclaves’ input and output data. The Runtime Attestor records a hash of each data item together with metadata that describes both the intended usage (e.g., input or output) and properties (e.g., model or dataset). Once an input/output pair is collected, it constructs the EDR as a nested JSON object: keys contain the usage and properties of the data, and the values contain data hashes. The Runtime Attestor invokes the secure processor’s signing API with a SHA-256 digest of the EDR. The secure processor signs it together with a system measurement, and returns the signed EDR.
6.3. Exclave emulation
To implement exclave attestation without relying on software secrets, we emulate the attestation path using a virtual Trusted Platform Module (vTPM) as the secure processor in AMD SEV-SNP.
The vTPM provides attestation primitives, including measurement and signing, and executes at the highest privilege level in the CVM. Its signing key is endorsed by the PSP when the CVM launches. We use virtual machine privilege modules (VMPL) to create privilege levels: the vTPM runs in VMPL0 (highest privilege), while the exclave runs in VMPL2 (lower privilege). The MMU and RMP enforce isolation so that the exclave in a lower privilege level cannot interfere with the attestation process at higher privilege.
Attestation then uses an MMU-protected shared memory page, which is a VMPL2 page writable by both the vTPM and the exclave. For attestation, the vTPM reads the runtime claim’s hash from the shared page, extends it into the PCR23, signs the PCRs, and writes the signed report back.
@@ 6.4. Secure storage with commitment
The Secure Storage module provides integrity-preserving and verifiable data storage for exclaves, which is used to protect both datasets and task code.
Datasets. VerifiableFL uses Linux’s dm-verity (dmverity) for integrity protection, which checks accessed disk blocks against a Merkle tree. Data providers use the dm-verity tool to package datasets with a Merkle tree. The generated data commitment becomes a unique measurement over the dataset, which is included in EDRs by the Runtime Attestor. When setting up an exclave, VerifiableFL mounts the data disk and checks the expected data commitment against the Merkle tree.
@@ 7. Evaluation
We evaluate VerifiableFL to answer the following questions: (i) what is VerifiableFL’s end-to-end training performance compared to existing solutions? (§7.2); (ii) how does VerifiableFL scale with an increased number of data providers and GPUs? (§7.3); (iii) what is the performance breakdown of VerifiableFL? (§7.4); (iv) what overhead does VerifiableFL’s runtime attestation add? (§7.5); (v) what is the impact of secure storage on throughput and latency? (§7.6); and (vi) what is the auditing overhead as VerifiableFL scales? (§7.7)
@@ 7.1. Experimental setup
Testbed. We deploy confidential VMs (CVMs) on Azure. CVMs without GPUs run on AMD EPYC 7763v CPUs; CVMs with GPUs run on AMD EPYC 9V84 CPUs, both with SEV-SNP and vTPM enabled. CVMs with GPUs have one NVIDIA H100 GPU with Hopper Confidential Computing extensions (h100cc) enabled.
On Azure, we deploy 5 CVMs (EC4as_v5) to execute 1 model update exclave and 4 NVFlare clients; 5 CVMs (EC8as_v5) to execute 4 DP exclaves and 1 NVFlare server; and 1 CVM (EC16as_v5) as the aggregation exclave. Training exclaves are deployed on 4 CVMs with CPUs (DC32as_v5) or 4 CVMs with GPUs (NCC40ads_H100_v5). Each training exclave mounts an Azure P30 data disk, which has up to 200 MB/s of throughput and support 5,000 IOPS.
VerifiableFL implementation. We implement VerifiableFL in C++ and Python. It uses PyTorch v2.5.0 with CUDA 12.4 for the FL tasks, Azure Guest Attestation Library v1.0.2 for CVM attestation, and NVFlare v2.2.1 (roth2022nvidia) as the FL framework. [^2] Our open-source artifacts will be released upon acceptance. As described in §6.3, we emulate the attestation behaviour of exclaves using a vTPM implementation (vtpm) running on the AMD SEV-SNP platform.
Baselines. We compare VerifiableFL with the two baselines: (i) NVFlare (No security) executes the FL workflow with the NVFlare server and clients in regular VMs. Each NVFlare node runs in a monolithic VM with SEV-SNP features disabled; and (ii) NVFlare (CVM) (guo2024trustworthy) hosts each NVFlare client, as well as the server, in its own monolithic CVM. For a fair performance comparison, both NVFlare (CVM) and VerifiableFL use the vTPM for attestation.
Datasets. We use 3 public datasets: (i) Wikitext103 (merity2016pointer) consists of raw text with over 100 M tokens from Wikipedia articles; (ii) NCBI-disease (dougan2014ncbi) is a disease text mining dataset with 793 fully annotated PubMed abstracts; and (iii) CIFAR10 (krizhevsky2009learning) has annotated $32\times 32$ pixel images with 50,000 training and 10,000 testing samples.
ML models. We use 6 widely-used ML models, covering three different categories: (i) decoder-only transformers—we use GPT-2 (radford2019language) with 124 M parameters, and a distilled version of GPT-2, DistilGPT2 (sanh2019distilbert), with 82 M parameters; (ii) encoder-only transformers—we employ BERT (kenton2019bert) with 110 M parameters, and MobileBERT (sun2020mobilebert) with 25 M parameters; and (iii) convolutional neural networks (CNN)—we use LeNet (lecun1998gradient), a 5-layer CNN, and VGG9 (simonyan2014very), a 9-layer CNN.
@@
Figure 8. Throughput with different numbers of clients
@@ 7.2. End-to-end throughput
We evaluate the end-to-end performance of VerifiableFL in terms of throughput. We run VerifiableFL with 1 FL server and 4 FL clients. Each client performs model training with either a CPU or GPU CVM, depending on the workload. After each training round, FL clients use the sparse vector differential privacy technique (SVT-DP) (lyu2016understanding) to add DP noise to the locally trained model. VerifiableFL runs each FL task for 10 rounds with federated averaging (FedAvg) aggregation. We measure the performance of VerifiableFL for our three ML categories: (i) language modeling; (ii) named-entity recognition; and (iii) image classification. Fig. 7 shows the measured end-to-end throughput results.
Language modeling. In this experiment, each client has a training exclave that executes the language modeling task in a GPU CVM (40 vCPUs; 320 GB RAM; 1 NVIDIA H100 GPU). We train both GPT-2 and DistilGPT2 models on the Wikitext dataset, which is partitioned into four subsets, one for each client.
@@
Fig. 7(a) shows the training throughput. With DistilGPT2, VerifiableFL has a $6\%$ overhead compared to the NVFlare (No security) baseline; NVFlare (CVM) has a $1\%$ overhead. With GPT-2, when compared to NVFlare (No security), VerifiableFL and NVFlare (CVM) incur less than $10\%$ and $4\%$ overhead in training throughput, respectively. We observe that, in both cases, VerifiableFL is slower than NVFlare (CVM): in VerifiableFL, exclaves introduce runtime overhead due to the runtime attestations; and the distributed deployment of exclaves introduces network overhead when (de)serializing the ML model and transmitting the serialized ML model.
Named-entity recognition. Training exclaves execute in CPU CVMs (32 vCPUs; 128 GB memory). We use the MobileBERT and BERT models on the NCBI-disease dataset, which is split into four shards as each client’s local dataset.
@@
Fig. 7(b) shows the training throughput. With MobileBERT, VerifiableFL and NVFlare (CVM) are $5\%$ and $4\%$ slower than NVFlare (No security), respectively. With BERT, VerifiableFL incurs a $9\%$ overhead compared to NVFlare (No security), and NVFlare (CVM) has a $4\%$ performance cost. We conclude that, when training MobileBERT, VerifiableFL has less performance impact compared to the BERT training. Since MobileBERT is $4.3\times$ smaller than BERT, it results in less network communication between exclaves in VerifiableFL.
Image classification. We also execute the image classification task with CPU CVMs and the same setup as before. Here, we use the LeNet and VGG9 models with the CIFAR10 dataset. Since the CIFAR10 dataset is relatively small, we replicate it four times and shard it into four subsets, ensuring each client receives sufficient training data.
Fig. 7(c) shows the performance results. For VGG9 training, VerifiableFL and NVFlare (CVM) incur $9\%$ and $4\%$ of overhead, respectively, when compared to NVFlare (No security). Compared to other workloads, VerifiableFL incurs a higher overhead for LeNet training: it is $11\%$ slower than NVFlare (No security). Even with NVFlare (CVM), LeNet training overhead is 7%. The significant CVM overhead with VerifiableFL’s extra scheduling and communication impact performance.
@@
For moderate and large models (i.e., VGG9, MobileBERT, BERT, DistilGPT2, and GPT-2), VerifiableFL’s overhead is below $10\%$ compared to NVFlare (No security), which is slightly higher than NVFlare (CVM)’s overhead of $4\%$. Across all models, the overhead stays low. LeNet has a slightly higher overhead than other models, because its small model requires little training time, which makes it sensitive to the CVM overhead and VerifiableFL’s runtime and I/O costs.